OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y - x = 5 (2 points) (-10, -5) (-7, -2) (7, 2) (5, 10)

OpenStudy (anonymous):

y^2 + x^2 = 53 y - x = 5 (-7,-2) check... -2^2 + (-7^2) = 53 4 + 49 = 53 53 = 53 (correct) y - x = 5 -2 - (-7) = 5 -2 + 7 = 5 5 = 5 (correct)